# Spatial4D-Bench: A Versatile 4D Spatial Intelligence Benchmark

### v1

@@ 1 Introduction
In cognitive science, spatial cognition seeks to understand how humans and animals perceive, interpret, mentally represent, and interact with the spatial characteristics of the environment [waller2013handbook, newcombe2004spatial]. To assess human spatial intelligence, a wide range of standardized tests have been developed over the past decades, such as puzzles, pattern blocks, tangrams, paper-and-pencil tests [carroll1993human, eliot1983international, hegarty2005individual], Mental Rotations Test (MRT) [peters1995redrawn], and navigation tests based on virtual 3D reality environments [wolbers2008spatial]. Analogously, Multimodal Large Language Models (MLLMs) have been developed to unify language and vision, with the ultimate goal of achieving human-level spatial understanding and reasoning. Recent developments in MLLMs [OpenAI2025, Gemini2025, zhang2025videollama, bai2025qwen2, wang2025internvl3] have achieved impressive performance, covering a wide range of multimodal understanding and reasoning tasks. However, to what extent MLLMs can achieve human spatial cognition levels remains an open question.
To investigate this, researchers have recently developed some spatial intelligence benchmarks [Yang_2025_CVPR, yu2025far, liu2025mirage, gong2025space, deng2025internspatial, yang2025mmsi, ma20243dsrbench, cheng2024spatialrgpt, zhao2025mmvu, tang2025lego] and test a variety of MLLMs to assess their spatial reasoning capabilities. These benchmarks predominantly assess foundational spatial reasoning capabilities, such as room size estimation, object distance reasoning, and object counting. Although these benchmarks have revealed substantial capability limitations of existing MLLMs on spatial reasoning tasks, they are often small-scale and lack diversity. This poses a challenge for comprehensively evaluating the spatial intelligence capabilities of MLLMs. Moreover, most existing spatial intelligence benchmarks focus only on 3D spatial intelligence with a primary focus on reasoning about static scenes, largely neglecting the evaluation of spatiotemporal awareness. Yet the real world is inherently a 4D environment, where spatial and temporal aspects continuously evolve. Living in such a time-evolving 4D environment, humans naturally develop 4D spatial intelligence in which perceiving and processing how objects move or change dynamically over time are the key tasks, enabling a variety of spatial reasoning abilities. In light of this, several spatial intelligence benchmarks have involved spatiotemporal reasoning, such as STI-Bench [li2025sti], VSI-SUPER [yang2025cambrian], and VLM4D [zhou2025vlm4d]. However, the evaluation data for these benchmarks remain limited in terms of diversity or scale, limiting their ability to comprehensively evaluate MLLMs in 4D spatial reasoning and to assess the gap between MLLMs and human-level 4D spatial intelligence.
@@
SpatialRGPT [cheng2024spatialrgpt] Jun. 2024 1,406 2
3DSRBench [ma20243dsrbench] Dec. 2024 2,772 4
VSI-Bench [Yang_2025_CVPR] Dec. 2024 5,156 8
STI-Bench [li2025sti] Mar. 2025 2,064 8
LEGO-Puzzles [tang2025lego] Mar. 2025 1,100 11
MMSI-Bench [yang2025mmsi] May. 2025 1,000 4
MIRAGE [liu2025mirage] May. 2025 1,710 3
SpaCE-10 [gong2025space] Jun. 2025 5,000 8
InternSpatial [deng2025internspatial] Jun. 2025 6,008 4
VLM4D [zhou2025vlm4d] Aug. 2025 1,816 4
SIBench [yu2025far] Sep. 2025 9,000 7
@@
In this work, we present Spatial4D-Bench, a large-scale, multi-task 4D spatial intelligence benchmark that enables comprehensive assessment of MLLMs’ spatial reasoning abilities. As shown in Figure˜1, Spatial4D-Bench comprises ~40,000 carefully curated and annotated question-answer (QA) pairs, covering a wide variety of indoor and outdoor environments involving diverse objects, actions, and scenes. By adhering to human spatial cognition principles [waller2013handbook, newcombe2004spatial], these QA pairs are divided into 6 categories, namely, object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning, and spatiotemporal reasoning. Each category is further subdivided into various tasks, yielding 18 tasks in total that span a broad range of spatial perception, understanding and reasoning abilities. This significantly distinguishes Spatial4D-Bench from existing benchmarks that are often small-scale or limited in diversity. Although existing benchmarks have covered some tasks (e.g., object size estimation and object counting) overlapping with Spatial4D-Bench, some 4D tasks presented in Spatial4D-Bench remain insufficiently investigated, including but not limited to spatial memory, state change detection, and physical plausibility reasoning. Therefore, compared to existing benchmarks, Spatial4D-Bench provides a more comprehensive evaluation suite for the assessment of MLLMs’ spatial cognition abilities, spanning a variety of tasks that parallel the versatility of human spatial intelligence. Table˜1 summarizes the statistical differences between Spatial4D-Bench and existing benchmarks.
@@
We conduct thorough experiments to benchmark a variety of state-of-the-art MLLMs on Spatial4D-Bench, including two proprietary MLLMs (GPT-5 [OpenAI2025] and Gemini 2.5-Pro [Gemini2025]) and several open-source MLLMs (VideoLLama3 [zhang2025videollama], Qwen2.5-VL [bai2025qwen2], Qwen3-VL [bai2025qwen2], and InternVL3.5 [wang2025internvl3]) with model sizes ranging from 7B to 241B parameters. Our extensive experiments and in-depth analysis reveal that MLLMs still exhibit a performance gap relative to humans in comprehensive 4D spatial reasoning. In particular, MLLMs have substantial limitations in a wide variety of 4D spatial reasoning aspects, such as route plan, egocentric reasoning, and physical plausibility reasoning. Nevertheless, we also observe that MLLMs surpass human performance on certain tasks, such as room size and object size estimation. This is reasonable, as humans generally struggle with tasks that require the accurate estimation of 3D scale in real world, while MLLMs can outperform humans by leveraging vast amounts of training data to provide prior knowledge. We hope that these findings provide valuable insights to the community and that the release of Spatial4D-Bench facilitates the development of more capable MLLMs toward human-level 4D spatial intelligence. More resources can be found on our project page: https://spatial4d-bench.github.io/spatial4d/.
@@ 2.1 Spatial Intelligence Benchmark
Multiple spatial intelligence benchmarks [Yang_2025_CVPR, yu2025far, liu2025mirage, gong2025space, deng2025internspatial, yang2025mmsi, ma20243dsrbench, cheng2024spatialrgpt, zhao2025mmvu, tang2025lego] have recently emerged to evaluate MLLM’s capabilities in spatial reasoning tasks. VSI-Bench utilizes [Yang_2025_CVPR] public 3D scene datasets including ScanNet [dai2017scannet], ScanNet++ [yeshwanth2023scannet++] and ARKitScenes [baruch2021arkitscenes] to construct 5,000 QA pairs that span eight 3D spatial cognition tasks, categorized into three types of configurational, measurement estimation and spatiotemporal reasoning. VLM4D [zhou2025vlm4d] uses both real-world clips and synthetic videos to produce over 2,000 high-quality QA pairs to explore models’ spatiotemporal abilities in translational and rotational motion, perspective shifts, motion continuity, and related dynamics. STI-Bench [li2025sti] is built from over 300 real-world videos with more than 2,000 QA pairs to test both static and dynamic spatial tasks. More recently, VSI-SUPER [yang2025cambrian] complements VIS-Bench by adding long videos in the benchmark to construct a large instruction-tuning dataset named VSI-590K. While these benchmarks emphasize various spatial capabilities in different scenarios, there is a lack of a comprehensive benchmark that unifies multi-task spatial understanding and reasoning in the context of a unified 4D representation. Unlike existing benchmarks, Spatial4D-Bench is a large-scale, multi-task evaluation benchmark comprising ~40,000 question-answer pairs covering 18 well-defined tasks which are systematically organized into 6 categories. Spatial4D-Bench also presents seven tasks that are important for spatial intelligence but have not been sufficiently investigated in current benchmarks. The statistical differences between Spatial4D-Bench and existing benchmarks are summarized in Table˜1.
@@ 2.2 Multimodal Large Language Models
Multimodal Large Language Models extend language models [gemini2023, llama2_2023] by integrating extra input modalities such as video, image, and audio, to enable advanced spatial capabilities including visual reasoning and scene understanding with temporal information [Alayrac2022Flamingo, chen2024internvl]. Qwen-VL [bai2025qwen2, qwen3] focuses on architectural design to improve high-resolution visual recognition and OCR/text understanding inside images. InternVL [wang2025internvl3, chen2024internvl] emphasizes scalable vision encoders and multimodal preference optimization to enable scalable training pipelines. LLaVA [liu2023improvedllava] introduces vision instruction tuning to teach the model to follow complex multimodal instructions and makes it possible to train high-quality vision language models with relatively little manually curated multimodal data. VideoLLama [zhang2025videollama] balances input flexibility and model efficiency to support long-video QA, temporal reasoning, temporal grounding. These improvements jointly impel MLLMs to be more capable of achieving a wide variety of challenging spatial tasks.
@@ 2.3 Spatial Reasoning with MLLMs
In addition to benchmarking MLLMs, some researchers have also proposed various techniques to enhance the spatial reasoning capabilities of MLLMs. SpatialRGPT [cheng2024spatialrgpt] develops automated data labeling pipelines to generate large quantities of 3D visual QA pairs for training, with the integration of depth modules and the alignment of depth information with 2D visual embeddings. Tang et al.  [tang2025sparkle] investigates whether enhancing basic spatial abilities such as direction comprehension, distance estimation, and localization can improve the overall spatial reasoning performance. Chen et al.  [chen2025spatial] reveals that successful spatial reasoning largely depends on the model’s ability to attend to task-relevant objects. Some other work [fan2025vlm, wu2025spatial, zhu2024llava] further lifts 2D encodings to 3D to recover implicit 3D structural information. While various efforts have been made to improve the spatial reasoning capabilities of MLLMs, thorough evaluation of the model performance on a comprehensive benchmark will provide insights and understanding of how the current limitations can be further addressed.
@@ 3.1 Overview
Spatial4D-Bench is a large-scale, multi-task 4D spatial intelligence benchmark designed to assess the 4D spatial intelligence of MLLMs. We name this benchmark “Spatial4D-Bench” as it is constructed from a large amount of video data grounded in 4D space and aims to evaluate the gap between MLLMs and human-level 4D spatial intelligence. Spatial4D-Bench consists of ~40,000 carefully curated and annotated QA pairs of 18 tasks, ranging from fundamental perception to complex reasoning. A detailed categorization of the 18 tasks is shown in Figure˜1.
To ensure a comprehensive evaluation, Spatial4D-Bench aggregates data from a wide variety of publicly available datasets, encompassing both indoor and outdoor environments. These datasets capture a rich variety of objects, actions, and scenes viewed from both egocentric and allocentric perspectives. Instead of merely aggregating existing labels, we generate novel QA pairs across multiple datasets to maximize task diversity. Crucially, adhering to human spatial cognition principles [waller2013handbook, newcombe2004spatial], we organize these 18 tasks into a hierarchical taxonomy composed of 6 core categories: object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning and spatiotemporal reasoning. Compared to existing benchmarks, Spatial4D-Bench provides a more comprehensive evaluation suite to assess the spatial cognitive abilities of MLLMs, spanning a variety of tasks that parallel the versatility of human spatial intelligence.
@@ 3.2 Benchmark Construction
Figure˜2 illustrates an overview of the pipeline used to construct Spatial4D-Bench. It can be seen that there are four stages in the construction pipeline, including data collection, data unification, QA pairs generation, and final human review.
@@
In the first stage, we collect a large amount of data from a variety of datasets, including Charades-Ego [sigurdsson2018charades], ARKitScenes [baruch2021arkitscenes], ScanNet [dai2017scannet], ScanNet++ [yeshwanth2023scannet++], 3RScan [wald2019rio], RoomTour3d [han2025roomtour3d], EPIC-KITCHENS [Damen2018EPICKITCHENS], ADL [youhome2022], EgoTaskQA [jia2022egotaskqa], HoloAssist [wang2023holoassist], IndustReal [schoonbeek2024industreal], YouCook2 [zhou2018towards], Video-MME [fu2024video], VideoPhy-2 [bansal2025videophy], and nuScenes [caesar2020nuscenes]. These datasets encompass various objects, actions, and scenes across indoor and outdoor scenarios, egocentric and allocentric perspectives, and both real and synthetic/generated data, contributing to the diversity and comprehensiveness of Spatial4D-Bench. We aggregate diverse data modalities from the participating datasets, including text, RGB videos, and point clouds (used only to facilitate QA annotation). Figure˜3 summarizes the source datasets that contribute to each task in Spatial4D-Bench. In addition, as shown in Figure˜2, this stage includes a human verification loop to filter out low-quality scans or incomplete annotations, ensuring that the input data meet the reliability requirements.
@@
In the second stage, we convert the collected data into a unified metadata structure following [Yang_2025_CVPR] since these data are collected from a variety of datasets and are heterogeneous. This standardization facilitates consistent downstream task generation and enables traceability for error correction during human review and verification. During this stage, we also preprocess image frames, sourced from RoomTour3D [han2025roomtour3d], for the route plan task by tagging room names on the relevant frames. This preprocessing disambiguates the route planning targets, ensuring that the model focuses on spatial planning capabilities rather than OCR capabilities. For action prediction and spatial memory tasks, we select and extract video clips with various lengths from EPIC-KITCHENS [Damen2018EPICKITCHENS], ADL [youhome2022], EgoTaskQA [jia2022egotaskqa], HoloAssist [wang2023holoassist], IndustReal [schoonbeek2024industreal], and Video-MME [fu2024video], to satisfy the particular design requiremets of the QA pairs, achieving consistency between a QA pair and its corresponding video clip in the context of the given tasks. As a result, a subset of videos from the above-mentioned datasets have been preprocessed in curating Spatial4D-Bench. Human review and verification are also used in this stage to ensure the correctness of data unification.
@@
In the final stage, experienced AI researchers perform the final human review to filter out ambiguous or incorrect QA pairs, e.g., checking whether a textual description uniquely identifies a target object. As illustrated in Figure˜2, errors detected at this stage will trigger a feedback loop where issues are traced back to the source metadata for correction. Finally, we select ~40,000 high-quality QA pairs to construct the Spatial4D-Bench. Some examples of QA pairs can be found in the appendix (Table˜6 and Table˜7).

### v2

@@ 1 Introduction
In cognitive science, spatial cognition seeks to understand how humans and animals perceive, interpret, mentally represent, and interact with the spatial characteristics of the environment [35, 25]. To assess human spatial intelligence, a wide range of standardized tests have been developed over the past decades, such as puzzles, pattern blocks, tangrams, paper-and-pencil tests [6, 14, 19], Mental Rotations Test (MRT) [28], and navigation tests based on virtual 3D reality environments [38]. Analogously, Multimodal Large Language Models (MLLMs) have been developed to unify language and vision, with the ultimate goal of achieving human-level spatial understanding and reasoning. Recent developments in MLLMs [26, 10, 47, 41, 36] have achieved impressive performance, covering a wide range of multimodal understanding and reasoning tasks. However, to what extent MLLMs can achieve human spatial cognition levels remains an open question.
To investigate this, researchers have recently developed some spatial intelligence benchmarks [42, 46, 22, 17, 13, 44, 24, 9, 48, 31] and test a variety of MLLMs to assess their spatial reasoning capabilities. These benchmarks predominantly assess foundational spatial reasoning capabilities, such as room size estimation, object distance reasoning, and object counting. Although these benchmarks have revealed substantial capability limitations of existing MLLMs on spatial reasoning tasks, they are often small-scale and lack diversity. This poses a challenge for comprehensively evaluating the spatial intelligence capabilities of MLLMs. Moreover, most existing spatial intelligence benchmarks focus only on 3D spatial intelligence with a primary focus on reasoning about static scenes, largely neglecting the evaluation of spatiotemporal awareness. Yet the real world is inherently a 4D environment, where spatial and temporal aspects continuously evolve. Living in such a time-evolving 4D environment, humans naturally develop 4D spatial intelligence in which perceiving and processing how objects move or change dynamically over time are the key tasks, enabling a variety of spatial reasoning abilities. In light of this, several spatial intelligence benchmarks have involved spatiotemporal reasoning, such as STI-Bench [21], VSI-SUPER [43], and VLM4D [50]. However, the evaluation data for these benchmarks remain limited in terms of diversity or scale, limiting their ability to comprehensively evaluate MLLMs in 4D spatial reasoning and to assess the gap between MLLMs and human-level 4D spatial intelligence.
@@
SpatialRGPT [9] Jun. 2024 1,406 2
3DSRBench [24] Dec. 2024 2,772 4
VSI-Bench [42] Dec. 2024 5,156 8
STI-Bench [21] Mar. 2025 2,064 8
LEGO-Puzzles [31] Mar. 2025 1,100 11
MMSI-Bench [44] May. 2025 1,000 4
MIRAGE [22] May. 2025 1,710 3
SpaCE-10 [17] Jun. 2025 5,000 8
InternSpatial [13] Jun. 2025 6,008 4
VLM4D [50] Aug. 2025 1,816 4
SIBench [46] Sep. 2025 9,000 7
@@
In this work, we present Spatial4D-Bench, a large-scale, multi-task 4D spatial intelligence benchmark that enables comprehensive assessment of MLLMs’ spatial reasoning abilities. As shown in Fig.˜1, Spatial4D-Bench comprises ~40,000 carefully curated and annotated question-answer (QA) pairs, covering a wide variety of indoor and outdoor environments involving diverse objects, actions, and scenes. By adhering to human spatial cognition principles [35, 25], these QA pairs are divided into 6 categories, namely, object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning, and spatiotemporal reasoning. Each category is further subdivided into various tasks, yielding 18 tasks in total that span a broad range of spatial perception, understanding and reasoning abilities. This significantly distinguishes Spatial4D-Bench from existing benchmarks that are often small-scale or limited in diversity. Although existing benchmarks have covered some tasks (e.g., object size estimation and object counting) overlapping with Spatial4D-Bench, some 4D tasks presented in Spatial4D-Bench remain insufficiently investigated, including but not limited to spatial memory, state change detection, and physical plausibility reasoning. Therefore, compared to existing benchmarks, Spatial4D-Bench provides a more comprehensive evaluation suite for the assessment of MLLMs’ spatial cognition abilities, spanning a variety of tasks that parallel the versatility of human spatial intelligence. Table˜1 summarizes the statistical differences between Spatial4D-Bench and existing benchmarks.
@@
We conduct thorough experiments to benchmark a variety of state-of-the-art MLLMs on Spatial4D-Bench, including two proprietary MLLMs (GPT-5 [26] and Gemini 2.5-Pro [10]) and several open-source MLLMs (VideoLLama3 [47], Qwen2.5-VL [41], Qwen3-VL [41], and InternVL3.5 [36]) with model sizes ranging from 7B to 241B parameters. Our extensive experiments and in-depth analysis reveal that MLLMs still exhibit a performance gap relative to humans in comprehensive 4D spatial reasoning. In particular, MLLMs have substantial limitations in a wide variety of 4D spatial reasoning aspects, such as route plan, egocentric reasoning, and physical plausibility reasoning. Nevertheless, we also observe that MLLMs surpass human performance on certain tasks, such as room size and object size estimation. This is reasonable, as humans generally struggle with tasks that require the accurate estimation of 3D scale in real world, while MLLMs can outperform humans by leveraging vast amounts of training data to provide prior knowledge. We hope that these findings provide valuable insights to the community and that the release of Spatial4D-Bench facilitates the development of more capable MLLMs toward human-level 4D spatial intelligence. More resources can be found on our project page: https://spatial4d-bench.github.io/spatial4d/.
@@ 2.1 Spatial Intelligence Benchmark
Multiple spatial intelligence benchmarks [42, 46, 22, 17, 13, 44, 24, 9, 48, 31] have recently emerged to evaluate MLLM’s capabilities in spatial reasoning tasks. VSI-Bench utilizes [42] public 3D scene datasets including ScanNet [11], ScanNet++ [45] and ARKitScenes [4] to construct 5,000 QA pairs that span eight 3D spatial cognition tasks, categorized into three types of configurational, measurement estimation and spatiotemporal reasoning. VLM4D [50] uses both real-world clips and synthetic videos to produce over 2,000 high-quality QA pairs to explore models’ spatiotemporal abilities in translational and rotational motion, perspective shifts, motion continuity, and related dynamics. STI-Bench [21] is built from over 300 real-world videos with more than 2,000 QA pairs to test both static and dynamic spatial tasks. More recently, VSI-SUPER [43] complements VIS-Bench by adding long videos in the benchmark to construct a large instruction-tuning dataset named VSI-590K. While these benchmarks emphasize various spatial capabilities in different scenarios, there is a lack of a comprehensive benchmark that unifies multi-task spatial understanding and reasoning in the context of a unified 4D representation. Unlike existing benchmarks, Spatial4D-Bench is a large-scale, multi-task evaluation benchmark comprising ~40,000 question-answer pairs covering 18 well-defined tasks which are systematically organized into 6 categories. Spatial4D-Bench also presents seven tasks that are important for spatial intelligence but have not been sufficiently investigated in current benchmarks. The statistical differences between Spatial4D-Bench and existing benchmarks are summarized in Table˜1.
@@ 2.2 Multimodal Large Language Models
Multimodal Large Language Models extend language models [33, 1] by integrating extra input modalities such as video, image, and audio, to enable advanced spatial capabilities including visual reasoning and scene understanding with temporal information [2, 8]. Qwen-VL [41, 40] focuses on architectural design to improve high-resolution visual recognition and OCR/text understanding inside images. InternVL [36, 8] emphasizes scalable vision encoders and multimodal preference optimization to enable scalable training pipelines. LLaVA [23] introduces vision instruction tuning to teach the model to follow complex multimodal instructions and makes it possible to train high-quality vision language models with relatively little manually curated multimodal data. VideoLLama [47] balances input flexibility and model efficiency to support long-video QA, temporal reasoning, temporal grounding. These improvements jointly impel MLLMs to be more capable of achieving a wide variety of challenging spatial tasks.
@@ 2.3 Spatial Reasoning with MLLMs
In addition to benchmarking MLLMs, some researchers have also proposed various techniques to enhance the spatial reasoning capabilities of MLLMs. SpatialRGPT [9] develops automated data labeling pipelines to generate large quantities of 3D visual QA pairs for training, with the integration of depth modules and the alignment of depth information with 2D visual embeddings. Tang et al.  [32] investigates whether enhancing basic spatial abilities such as direction comprehension, distance estimation, and localization can improve the overall spatial reasoning performance. Chen et al.  [7] reveals that successful spatial reasoning largely depends on the model’s ability to attend to task-relevant objects. Some other work [15, 39, 51] further lifts 2D encodings to 3D to recover implicit 3D structural information. While various efforts have been made to improve the spatial reasoning capabilities of MLLMs, thorough evaluation of the model performance on a comprehensive benchmark will provide insights and understanding of how the current limitations can be further addressed.
@@ 3.1 Overview
Spatial4D-Bench is a large-scale, multi-task 4D spatial intelligence benchmark designed to assess the 4D spatial intelligence of MLLMs. We name this benchmark “Spatial4D-Bench” as it is constructed from a large amount of video data grounded in 4D space and aims to evaluate the gap between MLLMs and human-level 4D spatial intelligence. Spatial4D-Bench consists of ~40,000 carefully curated and annotated QA pairs of 18 tasks, ranging from fundamental perception to complex reasoning. A detailed categorization of the 18 tasks is shown in Fig.˜1.
To ensure a comprehensive evaluation, Spatial4D-Bench aggregates data from a wide variety of publicly available datasets, encompassing both indoor and outdoor environments. These datasets capture a rich variety of objects, actions, and scenes viewed from both egocentric and allocentric perspectives. Instead of merely aggregating existing labels, we generate novel QA pairs across multiple datasets to maximize task diversity. Crucially, adhering to human spatial cognition principles [35, 25], we organize these 18 tasks into a hierarchical taxonomy composed of 6 core categories: object understanding, scene understanding, spatial relationship understanding, spatiotemporal relationship understanding, spatial reasoning and spatiotemporal reasoning. Compared to existing benchmarks, Spatial4D-Bench provides a more comprehensive evaluation suite to assess the spatial cognitive abilities of MLLMs, spanning a variety of tasks that parallel the versatility of human spatial intelligence.
@@ 3.2 Benchmark Construction
Fig.˜2 illustrates an overview of the pipeline used to construct Spatial4D-Bench. It can be seen that there are four stages in the construction pipeline, including data collection, data unification, QA pairs generation, and final human review.
@@
In the first stage, we collect a large amount of data from a variety of datasets, including Charades-Ego [30], ARKitScenes [4], ScanNet [11], ScanNet++ [45], 3RScan [34], RoomTour3d [18], EPIC-KITCHENS [12], ADL [27], EgoTaskQA [20], HoloAssist [37], IndustReal [29], YouCook2 [49], Video-MME [16], VideoPhy-2 [3], and nuScenes [5]. These datasets encompass various objects, actions, and scenes across indoor and outdoor scenarios, egocentric and allocentric perspectives, and both real and synthetic/generated data, contributing to the diversity and comprehensiveness of Spatial4D-Bench. We aggregate diverse data modalities from the participating datasets, including text, RGB videos, and point clouds (used only to facilitate QA annotation). Fig.˜3 summarizes the source datasets that contribute to each task in Spatial4D-Bench. In addition, as shown in Fig.˜2, this stage includes a human verification loop to filter out low-quality scans or incomplete annotations, ensuring that the input data meet the reliability requirements.
@@
In the second stage, we convert the collected data into a unified metadata structure following [42] since these data are collected from a variety of datasets and are heterogeneous. This standardization facilitates consistent downstream task generation and enables traceability for error correction during human review and verification. During this stage, we also preprocess image frames, sourced from RoomTour3D [18], for the route plan task by tagging room names on the relevant frames. This preprocessing disambiguates the route planning targets, ensuring that the model focuses on spatial planning capabilities rather than OCR capabilities. For action prediction and spatial memory tasks, we select and extract video clips with various lengths from EPIC-KITCHENS [12], ADL [27], EgoTaskQA [20], HoloAssist [37], IndustReal [29], and Video-MME [16], to satisfy the particular design requiremets of the QA pairs, achieving consistency between a QA pair and its corresponding video clip in the context of the given tasks. As a result, a subset of videos from the above-mentioned datasets have been preprocessed in curating Spatial4D-Bench. Human review and verification are also used in this stage to ensure the correctness of data unification.
@@
In the final stage, experienced AI researchers perform the final human review to filter out ambiguous or incorrect QA pairs, e.g., checking whether a textual description uniquely identifies a target object. As illustrated in Fig.˜2, errors detected at this stage will trigger a feedback loop where issues are traced back to the source metadata for correction. Finally, we select ~40,000 high-quality QA pairs to construct the Spatial4D-Bench. Some examples of QA pairs can be found in the appendix (Table˜6 and Table˜7).